determine whether the graphs of each the following equation is parellel, perpindicular, or neither and show your work.
x+y=2,y=x+5

Answer by nerdybill(7384)

Convert the two given equations into the "slope-intercept" form:
y = mx + b
where
m is the slope
b is the y-intercept
.
x+y=2
y = -x + 2
y = (-1)x + 2
.
y=x+5
y = (1)x + 5
.
Since the slopes are "negative reciprocals" the two lines are perpendicular.